OpenStudy (anonymous):

sin^-1(1) exact value in degrees

OpenStudy (ddcamp):

\[\sin^{-1} (1) = \theta → \sin(\theta) = 1\] Are you familiar with the unit circle? 

OpenStudy (anonymous):

Yes, but it's been a while.